# Pauli Stabilizer Models for Gapped Boundaries of Twisted Quantum Doubles and Applications to Composite Dimensional Codes

### v3

@@ Acknowledgements.
We thank Phil Lotshaw, Paul Kairys, Sohan Ghosh, and Tyler Ellison for useful discussions and comments. M.M., A.G., and E.D. are supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Research Program, Early Career Award under contract number ERKJ420. This research was performed during M.M.’s participation in the Graduate Research at ORNL (GRO) Educational Program.
Appendix A Domain Wall between DS-$D(\mathbb{Z}_{4})$
In this appendix, we give the derivation of the domain wall between the DS phase and the $\mathbb{Z}_{4}$ phase using Alg. 1. This proves the more heuristic method given in Ex. IV.5. Before starting the algorithm, we place the $D(\mathbb{Z}_{4})$ on top while keeping the DS at the bottom. Denote the anyons of the bottom layer by $\{1,s_{1},\overline{s}_{1},b_{1}\}$ and for the top layer by $\{1,e_{2},m_{2},\dots\}$. The top layer will be a mirrored version of the bottom layer as in Fig. 23. The Hamiltonian is a mirrored version of the $\mathbb{Z}_{4}$ Hamiltonian Eq. (51). Here, the mirrored nature of the top layer will be fully exploited. Importantly, the anyon $e_{2}m_{2}$ with twist $\mathcal{T}(e_{2}m_{2})=i$ in the unfolded picture is mapped to the anyon $e_{2}m_{2}^{-1}$ with $\mathcal{T}^{\text{op}}(e_{2}m_{2})=i$ in the folded version because the handedness changes with mirror reflection.
We then begin the algorithm. First, we choose region $R$ extending to infinity. The boundary $\partial R$ lives on the direct lattice in both layers layers Fig. 51. Second, we add the shortest ribbon operators creating anyons of the Lagrangian subgroup generated by $\mathcal{L}_{\mathbb{Z}_{4}-DS}=\{s_{1}e_{2}m_{2}^{-1},\overline{s}_{1}e_{2}m_{2},b_{1}e_{2}^{2},b_{1}m_{2}^{2},1_{1}e^{2}m_{2},\dots\}\bigr)$. We show a subset of the generators for illustration:
@@ Appendix A Domain Wall between DS-$D(\mathbb{Z}_{4})$
In the fourth step, we form new stabilizers that are products of the old ones and commute with the new stabilizers. These are given by products of the removed stars with their adjacent removed plaquettes along with the $F$ operators of layer $1$. In addition, the product of the plaquette squared of the two layers.
@@
In the fifth step, we erase condensed qudits. These are all the qudits in the bulk of $R$ or on its boundary $\partial R$. After condensing qudits, we also removed all stabilizers that only act on condensed qudits for clarity, Fig. 53.
The sixth step is to restrict the stabilizers at the boundary to their uncondensed support. This results in the stabilizers shown in Eq. (82). Finally, we unfold layer $2$ to be put on the right. This will result in the stabilizer of the bulk of layer 22 ($\mathbb{Z}_{4}$) to be the same again as the un-mirrored $D(\mathbb{Z}_{4})$ redoing the horizontal reflection. At each edge, now only one qudit lives, so we can drop the layers’ index unambiguously. This results in the following stabilizers:
@@
After stitching the lattice back, we find the domain wall between the DS and $\mathbb{Z}_{4}$ phases shown in Fig. 54. This proves the results using the more heuristic method used in Ex. IV.5 with direct condensation.
@@
Appendix B Inverse domain wall of $D(\mathbb{Z}_{4})$ $(e,m)\to(e^{-1},m^{-1})$}
In this appendix, we use Alg.1 along with the folding method to construct the inverse domain wall of $D(\mathbb{Z}_{4})$ code. More importantly, this exemplifies how finite domain walls ending with 0D defects can be constructed systematically.
@@ Appendix B Inverse domain wall of $D(\mathbb{Z}_{4})$ $(e,m)\to(e^{-1},m^{-1})$}
These operators have to be added in place of the old removed stabilizers everywhere. However, they are most important near the boundary, as they will create the domain wall. Thus, we do not need to figure out those terms except at the boundaries. Importantly, right at the 0D defect, the stars will need to satisfy both boundaries. This is the first time the order of the group will matter. So far, everything was applicable to all $D(\mathbb{Z}_{n})$. In $D(\mathbb{Z}_{4})$, the square of stars $A_{1}^{2}(v)A_{2}^{2}(v)$ will commute with both boundaries. Fifth, we erase condensed qudits in the bulk of the three regions along with their boundaries. In Fig. 57, the new plaquette and star operators are shown at the three boundaries, while condensed qudits are erased.
Figure 57: The configuration after the fourth and fifth step. In the fourth step, products of old stabilizers that commute with the new measured operators are added. We only added them near the boundary, as the rest of them will be condensed. After the fifth step, condensed qudits are erased.
Sixth, we need to restrict the operators to their uncondensed support. This results in the operators of the last column of Eq. (86). Since there are no more qudits to condense, we are done. We can unfold the two layers, dropping the layer index. The operators at the boundary will transform to:
@@
This is a valid finite domain wall mapping anyons to their anyons with a 0D defect at its two ends. We see that qudits on the vertical line have been condensed. If it is desirable to have lower-weight stabilizers, we can stitch the two sides together. This is formally done by imposing the $XX^{\dagger}$ or $XX$ two-edge operators at the Hilbert space level. This stitching will transform the two qudits into one. The action of any operator that acted on the qudits can now be mapped into this new degree of freedom. For the stabilizers right at the boundary, we have:
@@
The star operators that share an edge with the trivial wall will remain unchanged, while those sharing an edge with the inverse domain will result in new star operators $A^{\text{op}}(v)$. Right at the 0D defect, we have the order $2$ operator $X^{2}X^{2}$, which cannot stitch the two sides back. We will be left with four weight five operators denoted by $D^{(1)}$, $D^{(2)}$, $E^{(1)}$, and $E^{(2)}$.

### v4

@@ Acknowledgements.
We thank Shawn Cui, Birgit Kaufmann, Ralph Kaufmann, Phil Lotshaw, Paul Kairys, Sohan Ghosh, and Tyler Ellison for useful discussions and comments. We thank Sabre Kais for access to the Bell computing cluster at Purdue University. M.M., A.G., and E.D. are supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Research Program, Early Career Award under contract number ERKJ420. This research was performed during M.M.’s participation in the Graduate Research at ORNL (GRO) Educational Program. This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
Appendix A Construction of Domain Wall between DS-$D(\mathbb{Z}_{4})$
In this appendix, we give the derivation of the domain wall between the DS phase and the $\mathbb{Z}_{4}$ phase using Alg. 2. This proves the more heuristic method given in Ex. V.5. Before starting the algorithm, we place the $D(\mathbb{Z}_{4})$ on top while keeping the DS at the bottom. Denote the anyons of the bottom layer by $\{1,s_{1},\overline{s}_{1},b_{1}\}$ and for the top layer by $\{1,e_{2},m_{2},\dots\}$. The top layer will be a mirrored version of the bottom layer as in Fig. 27. The Hamiltonian is a mirrored version of the $\mathbb{Z}_{4}$ Hamiltonian Eq. (46). Here, the mirrored nature of the top layer will be fully exploited. Importantly, the anyon $e_{2}m_{2}$ with twist $\mathcal{T}(e_{2}m_{2})=i$ in the unfolded picture is mapped to the anyon $e_{2}m_{2}^{-1}$ with $\mathcal{T}^{\text{op}}(e_{2}m_{2})=i$ in the folded version because the handedness changes with mirror reflection.
We then begin the algorithm. First, we choose a region $R$ extending to infinity. The boundary $\partial R$ lives on the direct lattice in both layers, Fig. 54. Second, we add the shortest ribbon operators creating anyons of the Lagrangian subgroup generated by $\mathcal{L}_{\mathbb{Z}_{4}-DS}=\{s_{1}e_{2}m_{2}^{-1},\overline{s}_{1}e_{2}m_{2},b_{1}e_{2}^{2},b_{1}m_{2}^{2},1_{1}e^{2}m_{2},\dots\}\bigr)$. We show a subset of the generators for illustration:
@@ Appendix A Construction of Domain Wall between DS-$D(\mathbb{Z}_{4})$
In the fourth step, we form new stabilizers that are products of the old ones and commute with the new stabilizers. These are given by products of the removed stars with their adjacent removed plaquettes, along with the $F$ operators of layer $1$. In addition, the product of the plaquette squared of the two layers.
@@
In the fifth step, we erase trivialized qudits. These are all the qudits in the bulk of $R$ or on its boundary $\partial R$. After condensing qudits, we also removed all stabilizers that only act on trivialized qudits for clarity, Fig. 56.
The sixth step is to restrict the stabilizers at the boundary to their non-trivialized support. This results in the stabilizers shown in Eq. (75). Finally, we unfold layer $2$ to be put on the right. This will result in the stabilizer of the bulk of layer 22 ($\mathbb{Z}_{4}$) to be the same again as the un-mirrored $D(\mathbb{Z}_{4})$ redoing the horizontal reflection. At each edge, now only one qudit lives, so we can drop the layers’ index unambiguously. This results in the following stabilizers:
@@
After stitching the lattice back, we find the domain wall between the DS and $\mathbb{Z}_{4}$ phases shown in Fig. 57. This proves the results using the more heuristic method used in Ex. V.5 with direct condensation.
@@
Appendix B Example of a $0D$ Defect Construction
In this appendix, we use Alg.2 along with the folding method to construct the inverse domain wall of $D(\mathbb{Z}_{4})$ code ($(e,m)\to(e^{-1},m^{-1})$}). More importantly, this exemplifies how finite domain walls ending with 0D defects can be constructed systematically.
@@ Appendix B Example of a $0D$ Defect Construction
These operators have to be added in place of the old, removed stabilizers everywhere. However, they are most important near the boundary, as they will create the domain wall. Thus, we do not need to figure out those terms except at the boundaries. Importantly, right at the 0D defect, the stars will need to satisfy both boundaries. This is the first time the order of the group will matter. So far, everything was applicable to all $D(\mathbb{Z}_{n})$. In $D(\mathbb{Z}_{4})$, the square of stars $A_{1}^{2}(v)A_{2}^{2}(v)$ will commute with both boundaries. Fifth, we erase trivialized qudits in the bulk of the three regions along with their boundaries. In Fig. 60, the new plaquette and star operators are shown at the three boundaries, while trivialized qudits are erased.
Figure 60: The configuration after the fourth and fifth step. In the fourth step, products of old stabilizers that commute with the new measured operators are added. We only added them near the boundary, as the rest of them will be condensed. After the fifth step, trivialized qudits are erased.
Sixth, we need to restrict the operators to their non-trivialized support. This results in the operators of the last column of Eq. (79). Since there are no more qudits to condense, we are done. We can unfold the two layers, dropping the layer index. The operators at the boundary will transform to:
@@
This is a valid finite-domain wall mapping anyons to their anyons with a 0D defect at each end. We see that qudits on the vertical line have been condensed. If it is desirable to have lower-weight stabilizers, we can stitch the two sides together. This is formally done by imposing the $XX^{\dagger}$ or $XX$ two-edge operators at the Hilbert space level. This stitching will transform the two qudits into one. The action of any operator that acted on the qudits can now be mapped into this new degree of freedom. For the stabilizers right at the boundary, we have:
@@
The star operators that share an edge with the trivial wall will remain unchanged, while those sharing an edge with the inverse domain will result in new star operators $A^{\text{op}}(v)$. Right at the 0D defect, we have the order $2$ operator $X^{2}X^{2}$, which cannot stitch the two sides back. We will be left with four weight-five operators denoted by $D^{(1)}$, $D^{(2)}$, $E^{(1)}$, and $E^{(2)}$.
@@
Appendix C Belief Propagation
We will describe here the belief-propagation implementation in more detail. It follows the case for qubits closely [154, 145] generalized to the four-dimensional qudits with extra modifications due to the fact that $\mathbb{Z}_{4}$ is not a field. We operate in the usual setting where the problem is stated as a Tanner graph [172] where qudit and check nodes are denoted by the variables $v$ and $u$, respectively. The check node $u_{i}$ is connected to the data node $v_{j}$ with an edge if $H_{ij}\neq 0$. Furthermore, generalizing the qubit case, the edges are now weighted with exactly $H_{ij}$.
Given as input a syndrome vector $s$ of the same length as the number of code checks, the goal is to estimate the error vector $x$, which has the same length as the number of qudit nodes. For each check node $u_{i}$ that has $V(u_{i})$ neighboring data qudits, it naturally satisfies the equation $\sum_{j=1}^{V(u_{i})}H_{ij}x_{v_{j}}=s_{u_{i}}$. Here, the goal is to find the estimated error $x$ that satisfies all the constraints. This is just again the usual $H.x=s$, but recast in terms of the graph variables. The check-to-data messages used some sort of an XOR operation in the qubit case (that can be captured with a simple product operation since it was over $\mathbb{Z}_{2}$) [154]. This naturally generalizes to convolution for the $\mathbb{Z}_{4}$ case. The convolution of two vectors $f$ and $g$ given the syndrome $s$ is defined as:
$(f\oplus g)(s)=\sum_{t\in\mathbb{Z}_{4}}f(t)\cdot g(s-t\bmod 4)$ (83)
Here, we use the function notation for the entries of the vectors. For example, for a vector $g$, $g(0)$ is its first element. It will also be useful to define iterated convolution (usually over the data qudit neighbors of a check node ($V(u_{i})$):
$\bigotimes_{k\neq j}\phi_{k}=\phi_{1}\oplus\cdots\oplus\phi_{j-1}\oplus\phi_{j+1}\oplus\cdots\oplus\phi_{V\left(u_{i}\right)}.$
This naively scales like $O(4^{V(u_{i})})$. However, it can be computed using the forward-backward check-node update $O(16\times V(u_{i}))$ or fast Fourier transform $O(4\log(4)\times V(u_{i}))$ [54, 185, 56]. We will describe the convolution using the simpler naive approach, while numerically, we have used the forward-backward check-node update.
Let us now begin the algorithm. First, we take as an input the probability of physical errors $p$, assuming for simplicity that they are uniform and over a pure Pauli channel (e.g., pure X noise). The messages will be 4-dimensional vectors carrying the information for each power of $X$. The probability of failure for any physical nose is then $P(v)=(1-p,\frac{p}{3},\frac{p}{3},\frac{p}{3})$ assuming the noise model described in Sec. IV.2. The initial messages from the data qudits to the check nodes are then $m_{v\to u}(a)=P(v,a)$ and from check to data qudits is just $m_{u\to v}(a)=\frac{1}{4}$ since we do not yet have any preference for the estimated error.
In the next iteration, the check node $u_{i}$ projects the messages it has into contributions to the syndrome it sees from its neighbor data qudit $u_{i}$. This is captured in successive steps. We first calculate the probability that the qudit contributes $y$ value to the syndrome: $\phi_{j}(a)=\sum_{\begin{subarray}{c}y\in Z_{4}\\ H_{ij}\cdot y\equiv a\pmod{4}\end{subarray}}m_{v_{j}\rightarrow u_{i}}(y)$. The check nodes send their estimated errors to their neighboring data qudits by doing a convolution over all the neighbors except the receiver:
$m_{u_{i}\rightarrow v_{j}}(a)\propto\left(\bigotimes_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{V\left(u_{i}\right)}\phi_{k}\right)\left(s_{u_{i}}-H_{ij}\cdot a\bmod 4\right)$ (84)
The message is normalized such that $\sum_{b\in\mathbb{Z}_{4}}m_{u_{i}\rightarrow v_{j}}(b)=1$. After the data node receives its messages, it updates its belief to
$q_{v_{j}}(a)\propto P\left(v_{j},a\right)\prod_{u_{i}\in U\left(v_{j}\right)}m_{u_{i}\rightarrow v_{j}}(a).$ (85)
Here, $U(v_{j})$ denotes the check nodes neighboring the data node $v_{j}$. After that, the data node sends a ratio of its cumulative belief to the received belief from each check node:
$m_{v_{j}\rightarrow u_{i}}(a)=\frac{q_{v_{j}}(a)}{m_{u_{i}\rightarrow v_{j}}(a)}$ (86)
Optionally, we apply a damping scheme to improve numerical stability. We have found that this parameter is not very crucial near the thresholds. However, a non-zero value close to 1 was empirically found to help convergence.
$m_{v_{j}\rightarrow u_{i}}^{(t)}(a)=\alpha\cdot m_{v_{j}\rightarrow u_{i}}^{\text{new }}(a)+(1-\alpha)\cdot m_{v_{j}\rightarrow u_{i}}^{(t-1)}(a)$ (87)
Finally, we use the qudit’s belief to compute a hard decision estimate of the error: $\hat{x}_{v_{j}}=\arg\!\max(q_{v_{j}})$. This new vector takes the index with the highest probability for all data qudits. Finally, we check convergence $H.\hat{x}=s$ every few iterations to reduce computational cost. If the estimated error satisfies the syndrome, the algorithm halts; otherwise, it continues with new iterations. When the algorithm fails to converge after a preset maximum number of iterations $T_{\text{max}}$ is reached, it passes the soft decisions vectors $q$ to the OSD step described in Sec. IV.3.3. The algorithm is summarized in Alg. 3.
Algorithm 3 Qudit Belief Propagation Decoder
1:procedure Qudit-BP($H$, $s$, $p$, $\alpha$, $T_{\text{max}}$)
2:  Initialize priors $P(v,x)\leftarrow(1-p,p/3,p/3,p/3)$ for all data qudits $v$.
3:  Set initial messages: $m_{v\to u}(a)\leftarrow P(v,a)$ and $m_{u\to v}(a)\leftarrow 1/4$.
4:  for $t=1$ to $T_{\text{max}}$ do
5:   Check-to-Data: Project incoming messages:
[…]
7:   Convolve to find check messages:
[…]
9:   Normalize $m_{u_{i}\to v_{j}}(a)$.
10:   Data-to-Check: Compute qudit belief:
[…]
12:   Divide out check message:
[…]
14:   Update with damping:
[…]
16:   Hard Decision: Estimate error:
[…]
18:   if $H\hat{x}\equiv s\pmod{4}$ then
19:     return $\hat{x} \triangleright$ BP converged
20:   end if
21:  end for
22:  return Soft decisions $q$ $\triangleright$ Pass to OSD-0
23:end procedure
We note that for the depolarizing channel, one has vectors of length $16$ instead of $4$. We could also use $\log$ probabilities to avoid underflow. However, near the threshold, this is not an issue, and using raw probabilities is slightly faster. We finally note that during the preparation of this work, a recent paper appeared that shows promise in modifying the Tanner graph in order to avoid the OSD step altogether [86]. It will be interesting to see if an adaptation of this method to the qudit code could provide further improvements.